OpenStudy (anonymous):

3(x-1) - (x + 5) = 2

OpenStudy (anonymous):

i dont understand this question! can someone please help!!

OpenStudy (uri):

Solve the parentheses first 3(x-1) - (x + 5) = 2 3x-3-x-5=2 3x-x-3-5=2 2x-8=2 2x=2+8 2x=10 x=10/2 x=5